# Interpretable Causal Inference for Analyzing Wearable, Sensor, and Distributional Data

### 17 Dec 2023 v1

@@ 1 Introduction
We prove that ADD MALTS can consistently estimate treatment effects with complex, distributional data. Via simulation, we demonstrate that ADD MALTS can more accurately estimate conditional average treatment effects than competing methods; we also show how ADD MALTS adds trustworthiness in the causal pipeline by validating whether treated and control units are comparable in subregions of the covariate space. Finally, we re-analyze data studying the effectiveness of CGMs in managing health risks in patients with type 1 diabetes, finding important insights about the data and CGMs.
@@ 3.1 Theoretical Results
Let and assume If for some then for some monotonically increasing, zero-intercept function
@@
Let Assumption 1 hold. Let and where is the distance of the nearest neighbor of unit . And let be the barycenter of the KNN’s outcomes. Then,
@@
As evident in Theorem 1, as decreases and the right hand side will decrease to 0. Because we work on a bounded covariate space, is guaranteed to decrease and will go to 0. Therefore, we can consistently estimate conditional barycenters and conditional average treatment effects.
@@ 4.1 CATE Estimation
Our first experiment evaluates how well a variety of baselines and ADD MALTS can estimate CATEs. We consider four data generative processes (DGPs). In each DGP, we generate our distributional outcomes as truncated normal distributions (truncated at standard deviations from the mean). In the “Linear,” “Variance,” and “Complex” DGPs, we sample scalar covariates while the “Dist Cov” DGP has scalar covariates and one distributional covariate. A summary of the distributional covariate is used to generate the distributional outcome in “Dist Cov.” We evaluate each methods ability to estimate the CATE, , using the percent Integrated Relative Error: Section “CATE Estimation Experimental Details” of the supplement expands on our experimental setup.
@@ 6 Conclusion
While we discuss ADD MALTS’ utility in the context of continuous glucose monitoring data, it can be highly useful for analyzing a variety of other data. For example, distributional representations have been helpful for summarizing images (Oliva et al.,, 2013; Yang et al.,, 2020; Zhang et al.,, 2022) and for summarizing survey data from across geographies (Gunsilius,, 2023). However, these complex data may benefit from being represented as multidimensional distributions. Future directions should consider extending ADD MALTS to the setting of multidimensional distributional data. Lemma 1 offers an important first step in this direction as our proof of the consistency of the estimated conditional barycenter can easily be extended to multidimensional outcomes. Additionally, uncertainty quantification and variance estimation using distributional outcomes is a difficult and challenging issue that future research should explore. We propose preliminary insights in the supplementary material. Finally, wearable devices are becoming increasingly popular, leading to large datasets with hundreds of thousands of patients (Nazaret et al.,, 2022). To accommodate such data, future research could benefit from extensions of ADD MALTS that scale to larger datasets without sacrificing accuracy or interpretability.
@@ Appendix B CATE Estimation Consistency Proofs
Assumption 2.
Let and assume If for some then for some monotonically increasing, zero-intercept function 111Please note that this assumption is slightly different than the assumption in the main paper. In the main paper, the assumption requires that this smoothness condition needs to hold only with the 2-Wasserstein distance. This assumption requires the smoothness condition to hold for any -Wasserstein distance with
@@ B.1 Proof of Barycenter Consistency (Lemma 1)
Lemma.
Let Assumption 2 hold. Additionally, let represent the lower and upper bounds on the support of the space of outcome distributions, Specifically, for any cumulative distribution function for all and for all Lastly, let be the distance of the nearest neighbor. Then, for all
where is estimated as the barycenter of unit ’s KNN’s outcomes.
@@
We will prove that the following inequality holds for any p-Wasserstein distance, So, it will hold for the 2-Wasserstein distance as well.
Please note that we fix the covariates and treatment assignments for all units. The probability distribution of our inequality is with respect to the observed outcomes for each unit.
Let represent the conditional barycenter of the outcome at and treatment . Let represent the K nearest neighbors to unit in the estimation set with assigned treatment We demonstrate that ADD MALTS’ estimate – – consistently estimates the true conditional barycenter’s quantile function.
Now, let First, notice that by using a series of triangle inequalities, we can rewrite
Because we know that Then, by Assumption 2, we know that So, we can replace
Now, realize that is the centroid of the outcome distributions of the matched group. So, by definition,
because is either the centroid of the KNN set or is on average further away from any of the distributions in the set. Then, by using the triangle inequality once again, we can decompose this inequality once more:
where the last step follows from the smoothness assumption. Now, we can combine all of these small inequalities together:
Now, recall that so In this space, the maximum distance occurs when we move the entire distribution’s mass (i.e., 1) as far as possible, which is . So, Additionally, because is a distance metric, the -Wasserstein distance between distributions is bounded from below by 0. Because is bounded and each unit is independent, we can then bound this probability using Hoeffding’s inequality: for some
Let Then,
@@ B.2 Proof of CATE Consistency (Theorem 1)
Recall that the 1-Wasserstein distance between one-dimensional distributions is a contrast between quantile functions. So,
From Lemma 1, we know that
Therefore,
where is the maximum distance between unit and any of its treated or control K nearest neighbors.
Therefore,
@@ Appendix C Uncertainty Quantification
In this section, we provide different strategies for quantifying uncertainty around (i) sample average treatment effect on treated (SATT) and (ii) conditional average treatment effect estimates.

### 20 Mar 2024 v2

@@ 1 INTRODUCTION
We prove that ADD MALTS can consistently estimate treatment effects with complex, distributional data. Via simulation, we demonstrate that ADD MALTS can more accurately estimate conditional average treatment effects than competing methods; we also show how ADD MALTS adds trustworthiness in the causal pipeline by validating whether treated and control units are comparable in subregions of the covariate space. Finally, we re-analyze data studying the effectiveness of CGMs in managing health risks in patients with type 1 diabetes, finding important insights about the data and CGMs.111All code to replicate simulation and real data analysis results are available here: https://github.com/almost-matching-exactly/addmalts
@@ 3.1 Theoretical Results
Let and assume If for some then for some monotonically increasing, zero-intercept function where represents the Wasserstein distance.
Note that this assumption is under the -Wasserstein metric, which is measured as the largest pointwise distance between quantiles: The upper bounds all other -Wasserstein distances (Remark 6.6 in Villani et al., (2009)), so this assumption suggests that conditional barycenters become more similar with respect to any Wasserstein metric as the covariates become more similar.
@@
Let Assumption 1 hold. Let be the barycenter of the KNN’s outcomes. Assume the quantile functions of the distributions are Lipschitz continuous in the probability: there exists an such that for Let , where represents the distance from unit to its nearest neighbor according to . Then, there exists a function that is exponentially decreasing to 0 in such that
where is the 1-Wasserstein metric.
@@
As evident in Theorem 1, for any , the right hand side will decrease to 0 as Because we work on a bounded covariate space, is guaranteed to decrease and will go to 0 as the size of the estimation set increases. Therefore, we can consistently estimate conditional barycenters and conditional average treatment effects.
@@ 4.1 CATE Estimation
Our first experiment evaluates how well a variety of baselines and ADD MALTS can estimate CATEs. We consider four data generative processes (DGPs). In each DGP, we generate our distributional outcomes as truncated normal distributions (truncated at standard deviations from the mean). In the “Linear,” “Variance,” and “Complex” DGPs, we sample scalar covariates while the “Dist Cov” DGP has scalar covariates and one distributional covariate. A summary of the distributional covariate is used to generate the distributional outcome in “Dist Cov.”
We evaluate each methods ability to estimate the CATE, , using the percent Integrated Relative Error: Section “CATE Estimation Experimental Details” of the supplement expands on our experimental setup.
@@ 6 CONCLUSION
While we discuss ADD MALTS’ utility in the context of continuous glucose monitoring data, it can be highly useful for analyzing a variety of other data. For example, distributional representations have been helpful for summarizing images (Oliva et al.,, 2013; Yang et al.,, 2020; Zhang et al.,, 2022) and for summarizing survey data from across geographies (Gunsilius,, 2023). However, these complex data may benefit from being represented as multidimensional distributions. Future directions should consider extending ADD MALTS to the setting of multidimensional distributional data. Additionally, uncertainty quantification and variance estimation using distributional outcomes is a difficult and challenging issue that future research should explore. We propose preliminary insights in the supplementary material. Finally, wearable devices are becoming increasingly popular, leading to large datasets with hundreds of thousands of patients (Nazaret et al.,, 2022). To accommodate such data, future research could benefit from extensions of ADD MALTS that scale to larger datasets without sacrificing accuracy or interpretability.
ACKNOWLEDGEMENTS
We gratefully acknowledge support from NIH/NIDA R01DA054994, NIH/NIAID R01AI143381, DOE DE-SC0023194, NSF IIS-2147061, and NSF IIS-2130250. Alexander Volfovsky is also supported by a National Science Foundation Faculty Early Career Development Award (CAREER: Design and analysis of experiments for complex social processes). National Institute on Drug Abuse (NIDA) R01DA056407 supports Harsh Parikh. The authors want to thank Quinn Lanners for his insightful and constructive comments.
@@ Appendix B CATE Estimation Consistency Proofs
Assumption (1).
Let and assume If for some then for some monotonically increasing, zero-intercept function where represents the Wasserstein distance.
@@ B.1 Proof of Barycenter Consistency (Lemma 1)
Lemma (1).
Let Assumption 1 hold. Let be the barycenter of the KNN’s outcomes. Assume the quantile functions of the distributions are Lipschitz continuous in the probability: there exists an such that for Let , where represents the distance from unit to its nearest neighbor according to . Then, there exists a function that is exponentially decreasing to 0 in such that
where is the 1-Wasserstein metric.
@@
Let represent the conditional barycenter of the outcome at and treatment . Let represent the K nearest neighbors to unit in the estimation set with assigned treatment We demonstrate that ADD MALTS’ estimate – – consistently estimates the true conditional barycenter’s quantile function. Let
We begin by showing the following: for a given
Recall that the quantile function of the barycenter of one-dimensional, continuous distributions is the average of quantile functions, so
We manipulate the probability statement to bring it into the following form:
Recall that the -Wasserstein distance in one dimension is simply a contrast between quantile functions: Thus,
Recall that unit and unit are within distance of each other. By Assumption 1, we then know that So,
Then,
By Hoeffding’s inequality, we then know that
| (Pointwise Bound) |
This result demonstrates point-wise consistency at each probability level. Now, we will show consistency of estimating the conditional barycenter simultaneously.
Assume that all quantile functions, are Lipschitz continuous. Define the function
Because the average and difference of Lipschitz functions are Lipschitz, we then know that is also Lipschitz continuous. Let be the Lipschitz constant of
Let represent a ball of radius centered at . Choose an and a corresponding grid of points such that the following are satisfied: (1) and (2) and for any in Choose any Because all are within distance of , we know that Then,
Then,
| (Covering Bound) |
Now, let us return to our original probabilistic statement:
@@ B.2 Proof of CATE Consistency (Theorem 1)
where the last line follows from the definition of 1-Wasserstein in one-dimension.
Let be the maximum distance between unit and any of its treated or control K nearest neighbors. From Lemma 1, we know that
for an Therefore,